OpenStudy (anonymous):

4 is a solution of x < -5 Answer True or False

OpenStudy (anonymous):

Yes its true.

OpenStudy (anonymous):

can u explain why?

OpenStudy (anonymous):

The problem states the solutions of x are less than but not equal to -5. Is 4 less than -5?

OpenStudy (anonymous):

Yes of course because -5 is less than x so the x may be any number that is greater than -5 It would be -4,-3,-2,-1,0,1,2,3,4,...

OpenStudy (anonymous):

Rethink that rio96

OpenStudy (anonymous):

Try plugging 4 into x and see what you get.

OpenStudy (anonymous):

Sorry sorry it is false I got mistake in looking on problem.it is false

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do u have more questions then ask

OpenStudy (anonymous):

what about this one? -3 is a solution of y > -2 True or false?

OpenStudy (anonymous):

Use the logic I just posted. Try plugging -3 into y. Does the solution make sense?

OpenStudy (anonymous):

false?

OpenStudy (anonymous):

You pretty cute.

OpenStudy (anonymous):

Yes it is false. Because y is representing -3 less than -2 but in this question -3is greater than -2 so it is false.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but why is you fake?

OpenStudy (anonymous):

Do u have more questions? Then tag me there.

OpenStudy (anonymous):

Write the given equation in slope-intercept form, y = mx + b, by solving for y. 4x – 2y = 10 Answer y = 2x – 5 y = -4x – 5 y = 2x + 5 y = -2x – 5 @rio96

OpenStudy (anonymous):

The point (5, -1) is a solution of the equation: y = 2x − 11 True or false

OpenStudy (anonymous):

Can u write these question from new one